An organic compound was found to contain 54.5 % C, 9.2 % H, and 36.3 % O.



a. What is the empirical formula of the compound?



b. If the molar mass of the compound is 88.1 g/mol, what is its molecular formula?

1
Expert's answer
2022-03-12T02:01:03-0500

Assuming 100 g of the compound, the moles of each element are:


"n(C)=\frac{54.5\ g}{12.01\ g/mol}=4.54\ mol"


"n(H)=\frac{9.2\ g}{1.01\ g/mol}=9.11\ mol"


"n(O)=\frac{36.3\ g}{16.0\ g/mol}=2.27\ mol"


a) The lowest whole number ratio of C : H : O is 2 : 4 : 1. Therefore, the empirical formula is C2H4O.


b) The molar mass of 1 empirical formula unit is 44.05 g/mol. The molar mass of the compound is twice the molar mass of the empirical formula ("\frac{88.1}{44.05}=2" ). Therefore, the subscripts should be multiplied by 2 to get the molecular formula:

C4H8O2

Answer: a) C2H4O; b) C4H8O2